A physics class has 40 students. Of​ these, 15 students are physics majors and 17 students are female. Of the physics​ majors, six are female. Find the probability that a randomly selected student is female or a physics major.

The probability that a randomly selected student is female or a physics major is

nothing

.

​(Round to three decimal places as​ needed.)
Answer #1

Siolution:
General addition rule for any two events:
P(A or B) = P(A) + P(B) - P(A and B)

15 of the 40 students are physics majors = 15/40 = 0.375

17 of the 40 students are female = 17/40 = 0.425

06 of the 40 students are physics majors and female = 6/40 = 0.150

The probability that a randomly selected student is female or a physics major is 0.650
  
= 0.375 + 0.425 - 0.15 = 0.650